# A Pairwise Comparison Relation-assisted Multi-objective Evolutionary Neural Architecture Search Method with Multi-population Mechanism

### v1

@@ Abstract
Neural architecture search (NAS) enables researchers to automatically explore vast search spaces and find efficient neural networks. But NAS suffers from a key bottleneck, i.e., numerous architectures need to be evaluated during the search process, which requires a lot of computing resources and time. In order to improve the efficiency of NAS, a series of methods have been proposed to reduce the evaluation time of neural architectures. However, they are not efficient enough and still only focus on the accuracy of architectures. In addition to the classification accuracy, more efficient and smaller network architectures are required in real-world applications. To address the above problems, we propose the SMEM-NAS, a pairwise comparison relation-assisted multi-objective evolutionary algorithm based on a multi-population mechanism. In the SMEM-NAS, a surrogate model is constructed based on pairwise comparison relations to predict the accuracy ranking of architectures, rather than the absolute accuracy. Moreover, two populations cooperate with each other in the search process, i.e., a main population guides the evolution, while a vice population expands the diversity. Our method aims to provide high-performance models that take into account multiple optimization objectives. We conduct a series of experiments on the CIFAR-10, CIFAR-100 and ImageNet datasets to verify its effectiveness. With only a single GPU searching for 0.17 days, competitive architectures can be found by SMEM-NAS which achieves 78.91% accuracy with the MAdds of 570M on the ImageNet. This work makes a significant advance in the important field of NAS.footnotetext: This work has been submitted to the IEEE for possible publication. Copyright may be transferred without notice, after which this version may no longer be accessible.
@@ I Introduction
Convolutional neural networks (CNNs) have achieved great success in various machine learning tasks [1, 2, 3, 4]. However, in order to achieve promising performance, traditional CNNs are usually designed manually by human experts with extensive domain knowledge and experience. Not every interested user has such expertise, and even for experts, designing CNNs is also a time-consuming and error-prone process. Therefore, in order to facilitate and automate the design of deep convolutional neural networks, Zoph et al. first proposed the concept of neural architecture search (NAS) [5] and developed a well-known neural architecture search method for CNNs called NASNet [6], which can obtain neural architectures with highly competitive performance compared to hand-crafted neural architectures. In recent years, researchers have developed many NAS methods, which has attracted increasing attention from both industry and academia in a variety of learning tasks [7, 8], such as object detection [9], semantic segmentation [10], and natural language processing [11].
@@
In addition to the accuracy of neural networks, real-world applications also require NAS methods to find computationally efficient network architectures, e.g., low power consumption in mobile applications and low latency in autonomous driving. Generally speaking, the classification accuracy usually continuously increases with the complexity of the network architectures (i.e., the number of layers, the number of channels, etc.) [12]. This implies that maximizing the accuracy and minimizing the network complexity are two competing and conflicting objectives, thus requiring multi-objective optimization for NAS. Existing NAS algorithms can be classified into three categories including reinforcement learning (RL) based [6, 13], evolutionary algorithm (EA) based [14, 15] and gradient-based methods [16, 17]. Despite recent advances in RL-based and gradient-based NAS methods, they are still not easily applicable to multi-objective NAS. Gradient-based methods in which the search space is continuous rely on gradient descent to optimize the neural architectures, while other objectives, such as complexity, can not be optimized by the loss function. So, multiple objectives are not easy to be optimized and the diversity of architectures could be missed in gradient-based methods. RL-based methods are more costly and time-consuming than the other two methods. Evolutionary algorithm is a search method based on evolutionary principles that searches for an optimal solution through the evolution and selection [14]. When dealing with multi-objective problems in NAS field, evolutionary algorithms are more adaptable and flexible.
Regarding multi-objective NAS, researchers have proposed many methods for finding high-performance architectures that satisfy multiple metrics simultaneously [18]. For example, MnasNet [19] can be used to design a multi-objective neural architecture search approach that optimizes both accuracy and real-world latency on mobile devices. MixNet [20] can be used to design a new mixed depthwise convolution with accuracy and computation as the optimization metrics. But these methods have lower efficiency in optimizing architectures. In addition to the problem of search efficiency, there is a lack of improvement in the search strategy.
@@
During the search process, there is often a lack of diversity in the generated architecture, leading to premature convergence to suboptimal solutions. This phenomenon can be attributed to several factors, including the selection operation favouring one particular objective to retain individuals and the inadequate evolutionary operators. These constraints often hinder the exploration of the entire Pareto front, thereby limiting the quality of the obtained architecture [21]. For example, NSGANetV1 [15] tends to search for architectures within a small range around a certain FLOPs value, which may lead to the convergence of the local optimal solution. Therefore, how to maintain the diversity of the population without sacrificing performance is a challenging problem[22, 23]. There exist a number of diversity preservation approaches in existing multi-objective evolutionary algorithms, which usually select diverse architectures from the current generation and put them to the next generation to increase the diversity of the population during the search [21]. Inspired by these methods, which have not been well applied in NAS, we improve the selection operation and design a multi-population mechanism to generate more diverse architectures.
@@
Nevertheless, regardless of the search space and search strategy used, one bottleneck in this field is the need to evaluate a large number of network architectures during the search process, which requires huge computational resources and time [24]. Many researchers have proposed lots of methods to improve the efficiency of NAS algorithms, including weight sharing [13, 17], population memory [14], early stopping mechanism [25], and so on. But most of these methods still have to explicitly or implicitly select numerous architectures and then perform training to evaluate them. Recently, some researchers have established surrogate models to virtually evaluate network architectures, which can execute the environment selection with the predicted accuracy, significantly reducing the evaluation time [26]. However, one of the key challenges in obtaining reliable surrogate models is that obtaining training samples (pairs of architecture and accuracy) is computationally expensive in some extend. In existing methods, sample utilization of surrogate models is not efficient. For example, PRE-NAS [27] trains and evaluates a large number of architectures to be as training samples for the surrogate model. But, these training samples are without special data processing and the quantity is very limited. Moreover, we find that surrogate models are not necessary to produce reliable accuracy estimates (accuracy in the absolute sense), as long as the predicted results are consistent with the true performance ranking of architectures. In the evolutionary algorithm, the selection operation needs to be performed according to the fitness values of the individuals, and only the individuals with high fitness are selected for subsequent evolution. Therefore, it is more important to obtain the ranking of the candidate architectures than to predict the true accuracy.
@@
To address the above problems, we propose an efficient algorithm, called SMEM-NAS, which is a surrogate-assisted multi-objective evolutionary algorithm with a multi-population mechanism for NAS in this work. Firstly, we establish an efficient surrogate model based on pairwise comparison relations to obtain the accuracy ranking of candidate architectures. Then we conduct practical training on the selected top-ranked architectures. Furthermore, to consider other performance metrics and increase the diversity of the algorithm, we propose a multi-objective evolutionary algorithm based on a multi-population mechanism. Our approach aims to provide a set of high-performance architectures that take into account multiple optimization objectives. We validate the effectiveness of the proposed method on an image classification task with standard datasets, CIFAR-10, CIFAR-100 and ImageNet. The computational results show that our method outperforms most state-of-the-art NAS methods. This work intends to make the following contributions to advance the field of NAS:
@@
    It proposes a method to optimize both the accuracy and network complexity for NAS. We choose the number of multiply-add operations (MAdds) as the second objective. It can facilitate the ENAS to achieve high-quality architectures with low computational cost.
@@ II Related Work
In recent years, with deep neural networks widely deployed in different applications and computing environments, researches and explorations for NAS are particularly attractive. In the following, Section II-A and Section II-B present studies related to multi-objective NAS and surrogate models for NAS. The introduction and usage of the supernet are presented in Section II-C.
@@ II-A Multi-objective NAS
In the early methods of NAS, most studies only focus on the accuracy of neural networks. This often leads to that the searched networks suffer from excessive computation time and model size. On the other hand, the accuracy is rarely the only metric to be considered in real-world applications. It is often necessary to consider other additional and conflicting goals specific to the deployment scenario in the real world, such as model size, model computation, inference latency, and so on. Thus, some researchers have tried to take these secondary objectives into account while searching for architectures with high accuracy. For example, Lu et al. consider both classification accuracy and model computation, and then use the non-dominated sorting genetic algorithm II (NSGA-II) as the multi-objective optimization method [15, 28]. Besides, Xue et al. propose a multi-objective evolutionary algorithm with a probability stack for NAS, which considers accuracy and time consumption [29]. Wang et al. adopt a multi-objective particle swarm optimization algorithm to simultaneously optimize classification accuracy and FLOPs [30]. Du et al. further refine the multi-objective optimization process in NAS with the reference point-based environment selection [31]. Most of these existing approaches rely on the evaluation of architectures, and are not efficient enough and still time-consuming [32, 33]. There lacks the improvement of multi-objective evolutionary algorithms (MOEAs) in these methods, which tends to choose the existing multi-objective evolutionary algorithms such as NSGA-II [34]. There exists the phenomenon of small model traps during the search process [35]. Specifically, during the early search, smaller models with fewer parameters but higher test accuracy tend to dominate larger models that are less accurate but have the potential for higher accuracy [36]. Moreover, unsuitable evolution operators or conflicting objective characteristics may also lead to low diversity of the population and the imbalance between diversity and convergence, which may cause the search results to fall into the local optima [37].
@@ II-B Surrogate Models
The main computational bottleneck in NAS is the performance evaluation of architectures. In recent years, many surrogate-based methods have been proposed to accelerate the evaluation process in evolutionary algorithms [38]. The PNAS proposed by Liu et al. from Google uses long short-term memory (LSTM) as a surrogate model to replace the training of deep neural networks [39]. Afterwards, Luo et al. propose a neural network architecture optimization method, NAO, by using MLP as a surrogate model which works better than PNAS [16]. E2EPP [40] uses an offline surrogate model based on random forest (RF) to predict the performance directly from encodings of architectures. The online surrogate model is further improved in NSGANetV2 [34], which uses the architecture search to guide the construction of the accuracy predictor and significantly reduces the number of training samples. Guo et al. exploit a ranking loss function to learn a predictor model that predicts the relative score of architectures [41]. They can directly score and rank the performance of architectures to assist the evaluation process in NAS. However, accuracy predictors constructed and surrogate models adopted in most existing methods are simple and could not make full use of training samples. Inaccurate accuracy predictions also tend to lead to rank disorder of architecture performance, which affects the environment selection. Recently, Wang et al. propose a neural architecture search method based on particle swarm optimization, which uses support vector machine (SVM) as a surrogate model to conduct the preliminary exploration on the individual comparison relationship [42]. Despite the impressive progress, they search in a specific space constructed by themselves rather than in a general search space. That makes their method difficult to compare with other related methods and the method is not general enough.
@@ II-C SuperNet
To further improve the search efficiency of the proposed algorithm, we choose to adopt the weight sharing technique. At first, we need a supernet that includes all searchable architectures which are sub-networks. In this paper, the supernet we use is OnceForAll [43]. OnceForAll supports variations in four factors: depth, width, convolutional kernel size, and image resolution. In OnceForAll, an overall network is firstly trained by using a large amount of GPU resources, i.e., a maximal network is constructed by taking all the searched hyperparameters of the architectures (depth, width and kernel size) to their maximal values. Then the supernet is trained with the progressive shrinking algorithm [43], in which the largest network is fine-tuned to support sub-networks and they are added to the sampling space by sharing weights. The weights inherited from the trained supernet are used as a warm-up for the gradient descent algorithm during our architecture search.
Figure 1: Overall Framework of the multi-objective algorithm based on a multi-population mechanism (MP-MOEA) and the surrogate model assisting the search process (SMEM-NAS). (a) Offspring are generated with a multi-population mechanism. (b) The illustration of the evaluation assisted by the surrogate model. (c) The flowchart of the proposed algorithm.

### v2

@@ Abstract
Neural architecture search (NAS) enables researchers to automatically explore vast search spaces and find efficient neural networks. But NAS suffers from a key bottleneck, i.e., numerous architectures need to be evaluated during the search process, which requires a lot of computing resources and time. In order to improve the efficiency of NAS, a series of methods have been proposed to reduce the evaluation time of neural architectures. However, they are not efficient enough and still only focus on the accuracy of architectures. In addition to the classification accuracy, more efficient and smaller network architectures are required in real-world applications. To address the above problems, we propose the SMEM-NAS, a pairwise comparison relation-assisted multi-objective evolutionary algorithm based on a multi-population mechanism. In the SMEM-NAS, a surrogate model is constructed based on pairwise comparison relations to predict the accuracy ranking of architectures, rather than the absolute accuracy. Moreover, two populations cooperate with each other in the search process, i.e., a main population guides the evolution, while a vice population expands the diversity. Our method aims to provide high-performance models that take into account multiple optimization objectives. We conduct a series of experiments on the CIFAR-10, CIFAR-100 and ImageNet datasets to verify its effectiveness. With only a single GPU searching for 0.17 days, competitive architectures can be found by SMEM-NAS which achieves 78.91% accuracy with the MAdds of 570M on the ImageNet. This work makes a significant advance in the important field of NAS. Our code is publicly available at https://github.com/ccz-enas/SMEM-NAS.
@@ I Introduction
Convolutional neural networks (CNNs) have achieved great success in various machine learning tasks [1, 2, 3, 4]. The performance of neural networks is determined by structures and hyperparameters. Bayesian optimization methods like SMAC3 [5] are effective hyperparameter optimization methods. The structures of networks are also critical to performance. However, in order to achieve promising performance, traditional CNNs are usually designed manually by human experts with extensive domain knowledge and experience. Not every interested user has such expertise, and even for experts, designing CNNs is also a time-consuming and error-prone process. Therefore, in order to facilitate and automate the design of deep convolutional neural networks, Zoph et al. first proposed the concept of neural architecture search (NAS) [6] and developed a well-known neural architecture search method for CNNs called NASNet [7], which can obtain neural architectures with highly competitive performance compared to hand-crafted neural architectures. In recent years, researchers have developed many NAS methods, which have attracted increasing attention from both industry and academia in a variety of learning tasks [8, 9], such as object detection [10], semantic segmentation [11], and natural language processing [12].
@@
In addition to the accuracy of neural networks, real-world applications also require NAS methods to find computationally efficient network architectures, e.g., low power consumption in mobile applications and low latency in autonomous driving. Generally speaking, the classification accuracy usually continuously increases with the complexity of the network architectures (i.e., the number of layers, the number of channels, etc.) [13]. This implies that maximizing the accuracy and minimizing the network complexity are two competing and conflicting objectives, thus requiring multi-objective optimization for NAS. Existing NAS algorithms can be classified into three categories including reinforcement learning (RL) based [7, 14], evolutionary algorithm (EA) based [15, 16] and gradient-based methods [17, 18]. Despite recent advances in RL-based and gradient-based NAS methods, they are still not easily applicable to multi-objective NAS. In gradient-based methods, the search space is continuous rely on gradient descent to optimize the neural architectures, while other objectives, such as complexity, can not be optimized by the loss function. So, multiple objectives are not easy to be optimized and the diversity of architectures could be missed in gradient-based methods. RL-based methods are more costly and time-consuming than the other two methods. Evolutionary algorithm is a search method based on evolutionary principles that searches for an optimal solution through the evolution and selection [15]. When dealing with multi-objective problems in NAS field, evolutionary algorithms are more adaptable and flexible.
@@
Regarding multi-objective NAS, researchers have proposed many methods for finding high-performance architectures that satisfy multiple metrics simultaneously [19]. For example, Tan et al. designed a multi-objective neural architecture search approach that optimizes both accuracy and real-world latency on mobile devices [20]. Tan and Le designed a new mixed depthwise convolution with accuracy and computation as the optimization metrics [21]. But these methods have lower efficiency in optimizing architectures. In addition to the problem of search efficiency, there is a lack of improvement in the search strategy. During the search process, there is often a lack of diversity in the generated architecture, leading to premature convergence to suboptimal solutions. This phenomenon can be attributed to several factors, including the selection operation favouring one particular objective to retain individuals and inadequate evolutionary operators. These constraints often hinder the exploration of the entire pareto front, thereby limiting the quality of the obtained architecture [22]. For example, NSGANetV1 [16] tends to search for architectures within a small range around a certain FLOPs value, which may lead to the convergence of the local optimal solution. Therefore, how to maintain the diversity of the population without sacrificing performance is a challenging problem[23, 24]. Lu et al. proposed an efficient method for searching neural network architectures based on the Once-for-All supernet, which allows for rapid screening [25]. However, this method lacks further processing of the population, which prevents it from effectively exploring both objectives during the search process, thereby making it difficult to maintain good population diversity. There exist a number of diversity preservation approaches in existing multi-objective evolutionary algorithms, like multi-population strategies, which have not been well applied in NAS. Additionally, existing multi-population strategies often use static population selection strategies for the next generation. We consider that different search stages should focus on different search preferences, i.e., early strategies for exploitation and late strategies for exploration. Therefore, we improve the selection operation and propose a multi-population mechanism to generate more diverse architectures.
@@
Nevertheless, regardless of the search space and search strategy used, one bottleneck in this field is the need to evaluate a large number of network architectures during the search process, which requires huge computational resources and time [26]. Many researchers have proposed lots of methods to improve the efficiency of NAS algorithms, including weight sharing [14, 18], population memory [15], early stopping mechanism [27], and so on. But most of these methods still have to explicitly or implicitly select numerous architectures and then perform training to evaluate them. Recently, some researchers have established surrogate models to virtually evaluate network architectures, which can execute the environment selection with the predicted accuracy, significantly reducing the evaluation time [28]. However, one of the key challenges in obtaining reliable surrogate models is that obtaining training samples (pairs of architecture and accuracy) is computationally expensive in some extend. In existing methods, sample utilization of surrogate models is not efficient. For example, Peng et al. trained and evaluated a large number of architectures to be as training samples for the surrogate model in PRE-NAS [29]. In NSGANetV2, Lu et al. used a self-adaptive ensemble model to fit the accuracy of the architecture samples with the original encoding of the individuals [25]. But, these methods achieve training samples without special data processing and the quantity is very limited. Moreover, we find that surrogate models are not necessary to produce reliable accuracy estimates (accuracy in the absolute sense), as long as the predicted results are consistent with the true performance ranking of architectures. In the evolutionary algorithm, the selection operation needs to be performed according to the fitness values of the individuals, and only the individuals with high fitness are selected for subsequent evolution. Therefore, it is more important to obtain the ranking of the candidate architectures than to predict the true accuracy.
@@
To address the above problems, we propose an efficient algorithm, called SMEM-NAS, which is a surrogate-assisted multi-objective evolutionary algorithm with a multi-population mechanism for NAS in this work. The main framework of the proposed algorithm adopts NSGANetV2. The training accuracy of the network and the number of multiply-add operations (MAdds) are selected as two optimization objectives. We propose corresponding measures to address the issues of insufficient population diversity and inadequate predictive capability of the surrogate model. Firstly, we establish an efficient surrogate model based on pairwise comparison relations to obtain the accuracy ranking of candidate architectures. Then we conduct practical training on the selected top-ranked architectures. Furthermore, to simultaneously optimize the accuracy and network complexity, we design a multi-objective evolutionary algorithm with a multi-population mechanism and a comparison relations-based surrogate model, improving both the efficiency of the search process and the diversity of the resulting architectures. Our approach aims to provide a set of high-performance architectures that take into account multiple optimization objectives. This work intends to make the following contributions to advance the field of NAS:
@@
    We validate the effectiveness of SMEM-NAS on the image classification task with standard datasets, CIFAR-10, CIFAR-100 and ImageNet. The proposed method reduces search time to 0.17 GPU days on the ImageNet. The search architecture accuracy outperforms existing methods at the same amount of computation (MAdds).
@@ II Related Work
In recent years, deep neural networks have been widely deployed in different applications and computing environments. Researches and explorations for NAS are particularly attractive. In the following, Section II-A and Section II-B present studies related to multi-objective NAS and surrogate models for NAS. The introduction and usage of the supernet are presented in Section II-C.
@@ II-A Multi-objective NAS
Evolutionary algorithms are widely utilized to tackle multi-objective optimization problems due to their effectiveness in exploring large and complex solution spaces. By simulating the biological evolution, evolutionary algorithms are more flexible to handle multi-objective problems [30]. With the growing demand for efficient models in deep learning, the idea of multi-objective optimization has been introduced to neural architecture search. In the early researches of NAS, most studies only focus on the accuracy of neural networks. This often leads to that the searched networks suffer from excessive computation time and model size. However, the accuracy is rarely the only metric to be considered in real-world applications. It is often necessary to consider other additional and conflicting goals specific to the deployment scenario in the real world, such as model size, model computation, inference latency, and so on. Thus, some researchers have tried to take these secondary objectives into account while searching for architectures with high accuracy. For example, Lu et al. considered both classification accuracy and model computation, and then use the non-dominated sorting genetic algorithm II (NSGA-II) as the multi-objective optimization method [16, 31]. Besides, Xue et al. proposed a multi-objective evolutionary algorithm with a probability stack for NAS, which considers accuracy and time consumption [32]. Wang et al. adopted a multi-objective particle swarm optimization algorithm to simultaneously optimize classification accuracy and FLOPs [33]. Du et al. further refined the multi-objective optimization process in NAS with the reference point-based environment selection [34]. Most of these existing approaches rely on the evaluation of architectures, and are not efficient enough and still time-consuming [35, 36]. Most studies do not improve the multi-objective evolutionary algorithms (MOEAs), which tends to just use the existing multi-objective evolutionary algorithms such as NSGA-II [25]. There exists the phenomenon of small model traps during the search process [37]. Specifically, during the early search, smaller models with fewer parameters but higher test accuracy tend to dominate larger models that are less accurate but have the potential for higher accuracy [38]. Moreover, unsuitable evolution operators or conflicting objective characteristics may also lead to low diversity of the population and the imbalance between diversity and convergence, which may cause the search results to fall into the local optima. The multi-population mechanism is an effective strategy for addressing multi-objective optimization problems [39]. By maintaining multiple populations, this technique can explore different regions in the search space, solve different optimization tasks simultaneously and so on. Recently, this mechanism has also been adopted in multi-objective NAS [40, 41]. However, existing methods often use multiple populations for different datasets or tasks. For example, Song et al.  [42] leverage the classification strengths of various population architectures in their preferred categories to enhance classification accuracy. These multi-population NAS methods do not focus on the issue of diversity. Our research goal is to achieve a balance between accuracy and efficiency in a single task, further enhancing search efficiency and result diversity.
@@ II-B Surrogate Models
The main computational bottleneck in NAS is the performance evaluation of architectures. In recent years, many surrogate-based methods have been proposed to accelerate the evaluation process in ENAS [43]. Liu et al. from Google used the long short-term memory (LSTM) as a surrogate model to replace the training of deep neural networks [44]. Afterwards, Luo et al. proposed a neural network architecture optimization method, NAO, by using MLP as a surrogate model which works better than PNAS [17]. Sun et al. used an offline surrogate model based on random forest (RF) to predict the performance directly for the encoded architectures [45]. Guo et al. exploited a ranking loss function to learn a predictor model that predicts the relative score of architectures [46]. They can directly score and rank the performance of architectures to assist the evaluation process in NAS. However, most existing surrogate models, including those designed as accuracy predictors, are simplistic and fail to fully utilize training samples. For instance, accuracy predictors, which directly estimate the accuracy of architectures, often suffer from low generalization due to limited training data. This inaccuracy can disrupt the predicted rankings (i.e., rank disorder), thereby compromising the reliability of environment selection during the search process. Recently, Wang et al. proposed a neural architecture search method based on particle swarm optimization, which uses support vector machine (SVM) as a surrogate model to conduct the preliminary exploration on the individual comparison relationship [47]. Despite the impressive progress, they search in a specific space constructed by themselves rather than in a general search space. This makes their method difficult to compare with related methods, and limits its generalizability.
@@ II-C SuperNet
Supernet is a powerful technique within NAS. The supernet is a hyper-set of all the candidate networks in the search space, which means that each candidate network can be regarded as being formed by some modules or paths from the supernet. With such technique, there are two mainly used methods to efficiently search. One of the methods uses progressive training during the search. DARTS [18] updates the architecture parameters by gradients and it can also be reflected in the supernet. Yang et al. [37] used NSGA-III to search for architectures and only updates the parameters in the subnet during training of every step. This method can efficiently search for potential candidates without a pre-trained supernet. Another kind of methods is the one-shot method, which relies on a pre-trained supernet. Guo et al.  [48] proposed the single path one-shot neural architecture search, SPOS. They trained a supernet, and then sampled and evaluated each sub-network. The emergence of FairNAS [49] and GreedyNAS [50] further expanded the search methods based on the supernets. OnceForAll [51] is another widely used method for one-shot NAS. It supports variations in four factors: depth, width, convolutional kernel size, and image resolution. The OnceForAll supernet is firstly trained by the progressive shrinking algorithm, in which the largest network is fine-tuned to support sub-networks and they are added to the sampling space by sharing weights. The weights inherited from the trained supernet are used as a warm-up for the gradient descent algorithm during our architecture search.
Figure 1: Overall framework of the multi-objective algorithm based on a multi-population mechanism and the surrogate model assisting the search process (SMEM-NAS). (a) Offspring are generated with a multi-population mechanism. (b) The illustration of the evaluation assisted by the surrogate model. (c) The flowchart of the proposed algorithm. A set of elite architectures are retained in each generation and are added into the archive $\mathcal{A}$, which contains architectures for evolution.